A natural gas was containing mixture of methane
and ethane only. On complete combustion of
10 litres of gas at STP, the heat evolved
was 474.6 kJ. Assuming DHcombCH4(g)
= –894 kJ/mol and DHcombC2H6(g) = –1500 kJ/mol,
the percentage of CH4 and C2H6 will be
respectively

Grade:11

1 Answers

Sonu Dulat
6 years ago
Let ch4 =x such that C2h6=10-x ltr respSo. Moles of ch4=x/22.4; moles of C2h6=10-x/22.4ATQ(-894*x/22.4)-(1500*10-x/22.4)=-474.6After solving moles of ch4 (x)=7.23 so %=72.3% of C2h6=27.7